Use the data in the following​ table, which lists​ drive-thru order accuracy at popular fast food chains. Assume that orders are randomly selected from those included in the table.

​Drive-thru Restaurant   

A

B

C

D

Order Accurate

314314

274274

243243

144144

Order Not Accurate

3131

5353

3636

2020

If one order is​ selected, find the probability of getting food that is not from Restaurant A.

The probability of getting food that is not from Restaurant A is

nothing.

​(Round to three decimal places as​ needed.)
Answer #1
Drive thru Restaurant
A B C D Total
Order Accurate 314 274 243 144 975
Order Not Accurate 31 53 36 20 140
Total 345 327 279 164 1115

The total number of people who made the order = 1115

The number of people who made order from Restaurant A = 314 + 31 = 345

Probability of getting food from Restaurant A= 345 / 1115 = 0.309

Now

The probability of getting food that is not from Restaurant A

= 1 - Probability of getting food from Restaurant A

= 1 - 0.309

= 0.691

The probability of getting food that is not from Restaurant A is 0.691.